# Quantum Zeno Effect in Noisy Integrable Quantum Circuits for Impurity Models

### v1

@@ III Noiseless XX-impurity Quench
$\frac{(J^{2}-1)e^{ik}-e^{-ik}}{(J^{2}-1)e^{-ik}-e^{ik}}=e^{i2(L+1)k}.$ (S.33)
There exists a bound mode with $\Im k>0$, described by $B=0$ such that the wave function is normalizable in the thermodynamic limit, i.e. $(J^{2}-1)e^{ik}-e^{-ik}=0$ which gives $k_{BM}=\frac{i}{2}\log(J^{2}-1)$ with bound mode energy $E=2\cos k=\frac{J^{2}}{\sqrt{J^{2}-1}}$ and $\Im k>0$ gives $J>\sqrt{2}$. Likewise, $A=0$ gives $k_{BM}=-\frac{i}{2}\log(J^{2}-1)$. After studying the spectrum and equilibrium physics, now we can study the quench dynamics of this problem. Since the model is non-interacting, we need to solve only the single-mode evolution and take the algebraic sum of each mode at the end of the calculation. We prepare the initial state as $F(j,t=0)=F_{0}(j)=\delta_{j,j_{0}}$, and we solve for $F(j,t)$. The dynamics of a wave function is given by the following equation
$\begin{split}F(x,t)&=\sum_{\ell,k}F^{*}_{k}(\ell)F_{0}(\ell)F_{k}(j)e^{i\epsilon_{k}t}\\ &=\sum_{k}F^{*}_{k}(j_{0})F_{k}(j)e^{i\epsilon_{k}t}.\end{split}$ (S.34)
Let us look at the simplest case where we prepare the fermion (spin flip) at the impurity site initial $j_{0}=0$ and ask how the fermion number (magnetization) changes as time evolves, In the Kondo phase $J<\sqrt{2}$
$n_{0}(t)=F(0,t)^{2}=\left(\int\frac{dk}{\pi}\frac{J^{2}\sin^{2}(k)}{J^{4}-4(J^{2}-1)\cos^{2}(k)}e^{-i2t\cos k}\right)^{2},$ (S.35)
@@
where the Kondo time $t_{K}=\frac{\sqrt{|J^{2}-1|}}{J^{2}}$ is the only scale in this integral. We verify this result numerically as shown in Fig.S.2. In the Bound Mode Phase, the same quantity can be expressed as the following
@@
Figure S.2: Dynamics of the spin at the impurity site with initial Néel state. The system is in bound mode phase $J=3$ (left panel) where the analytic curve of the time-dependent spin profile is given by Eq. (S.40) and numerical cure is calculated for $L=40$ with exact diagonalization. The system is in the Kondo phase with $J=0.5$ (right panel) where the analytic curve of time-dependent spin profile is given by Eq. (S.40) and numerical cure is calculated for $L=40$ with exact diagonalization.
$F_{K}(j,\ell,t)=\int_{0}^{\pi}\frac{dk}{2\pi}\frac{1}{J_{\ell}J_{j}}(e^{-ikj}+S_{b}(k)e^{ikj})(e^{ik\ell}+S_{b}(k)^{*}e^{-ik\ell})e^{-i2t\cos k}.$ (S.38)
In the bound mode phase,
$F_{B}(j,\ell,t)=F_{K}(j,\ell,t)+\frac{J^{2}}{J_{j}J_{\ell}}\frac{J^{2}-2}{J^{2}-1}\left[\frac{e^{-iE_{B}t}}{(\sqrt{J^{2}-1})^{j+\ell}}+\frac{e^{iE_{B}t}}{(-\sqrt{J^{2}-1})^{j+\ell}}\right].$ (S.39)
$S^{z}(\ell,t)=2\sum_{j\in\{j_{x}\}}|F(j,\ell,t)|^{2}-1.$ (S.40)
with $F=F_{K}$ or $F=F_{B}$ depending on which phase the system is in.
IV Liouvillian Spectrum
In the previous section we solved XX-impurity Model quenched from an initial state with only some spins flipped, i.e. |$\ket{\psi_{0}}=\prod_{j_{x}}\sigma^{+}_{j_{x}}|\downarrow,\downarrow,...\downarrow\rangle$. Now we want to ask the question of how noise affects this model. Using Eq. (S.29), we need to study the two-particle sector of this model to solve for the dynamics of the two-point correlation function.
In the Liouville-Fock space, our initial state (initial condition of the correlation matrix) with spin-flip at site $j_{x}$ corresponds to the initial density matrix |$\ket{\rho_{0}}\rangle=c^{\dagger}_{j_{x}\uparrow}c^{\dagger}_{j_{x}\downarrow}\ket{0}\rangle$. If we initially flip more spins, we just need to superpose them, i.e. |$\ket{\rho_{0}}\rangle=\sum_{x_{j}}c^{\dagger}_{x_{j}\uparrow}c^{\dagger}_{x_{j}\downarrow}\ket{0}\rangle$. Due to the $U(1)\times U(1)$ symmetry of the Liouvillian operator, the number of fermions and total magnetization are conserved separately. Thus, we only need to diagonalize the eigenvalue problem in Eq. (S.28) in two particles $N=2$ and one spin-flip $M=1$ sector.
The time-independent Schrödinger equation read from Eq.(S.29) is
$EG(x,y)=\left[J_{x}\Delta_{x}+J_{y}\Delta_{y}+4i\gamma\delta_{x,y}+2i\gamma(\delta_{x}+\delta_{y}-2)\right]G(x,y),$ (S.41)
@@ IV Liouvillian Spectrum
$G_{k,q}(x,y)=\frac{1}{\mathcal{N}}\mathcal{S}\sum_{\alpha_{1},\alpha_{2}=\pm 1}e^{i\alpha_{1}kx+i\alpha_{2}qy}\left[A(\alpha_{1}k,\alpha_{2}q)\theta(x-y)+B(\alpha_{2}q,\alpha_{1}k)\theta(y-x)\right],$ (S.42)
where $\mathcal{N}$ is the normalization factor such that $1=\sum_{x,y}|G_{k,q}(x,y)|^{2}$ and $\mathcal{S}$ is the symmetrizer. Note if the transformation in Eq.(S.26) is not performed, then there is a factor $(-1)^{y}$ multiplying the wavefunction shown in the main text. The amplitudes are related as follows:
$\begin{split}A(-k,q)&=K(k)A(k,q),\\ A(k,-q)&=K(q)A(k,q),\\ A(k,q)&=S(k,q)B(k,q),\end{split}$ (S.43)
@@
$K(k)=-\frac{e^{ik}+e^{-ik}-2i\gamma-J^{2}e^{ik}}{e^{ik}+e^{-ik}-2i\gamma-J^{2}e^{-ik}},$ (S.44)
@@
$S(k,q)=\frac{\sin k-\sin q+2\gamma}{\sin k-\sin q-2\gamma}.$ (S.45)
@@
$e^{2ik(L+1)}$ $=K(k)S(k,q)S(-k,q)$ (S.46)
$e^{2iq(L+1)}$ $=K(q)S(q,k)S(-q,k),$ (S.47)
which when written explicitly takes the form
$\begin{split}-\frac{2\cos k-2i\gamma-J^{2}e^{ik}}{2\cos k-2i\gamma-J^{2}e^{-ik}}\left(\frac{\sin k-\sin q+2\gamma}{\sin k-\sin q-2\gamma}\right)\left(\frac{\sin k+\sin q-2\gamma}{\sin k+\sin q+2\gamma}\right)&=e^{i2(L+1)k}\\ -\frac{2\cos q-2i\gamma-J^{2}e^{iq}}{2\cos q-2i\gamma-J^{2}e^{-iq}}\left(\frac{\sin k-\sin q-2\gamma}{\sin k-\sin q+2\gamma}\right)\left(\frac{\sin k+\sin q-2\gamma}{\sin k+\sin q+2\gamma}\right)&=e^{i2(L+1)q}.\end{split}$ (S.48)
@@
IV.1 1. The free band (Roots with vanishing Imaginary part)
By taking the norm of Eq. (S.48), it is easy to see that
$k^{I}=-\frac{1}{2(L+1)}\ln\lvert S_{b}(k)S(k,q)S(-k,q)\rvert,$ (S.49)
@@
IV.2 Roots with Finite Imaginary Part
As mentioned earlier, when the scattering matrix is singular, the Bethe roots have a non-vanishing imaginary part in the thermodynamic limit. Now let us consider different situations with different parts of the S-matrix being singular.
Figure S.3: The upper two panels show the Liouvillian spectrum (top-left) and amplitude for the wavefunction in the double-bound mode (top-right) for $J=3$, $\gamma=0.5$ and $L=30$. The perturbative calculation matches the exact diagonalization result better since $\gamma$ is small. The lower two panels show the Liouvillian spectrum (bottom-left) and amplitude for the wavefunction in the double-bound mode (bottom-right) for $J=3$, $\gamma=2$ and $L=30$. The two-site Hamiltonian calculation matches the exact diagonalization result better since $\gamma$ is big. Also, wavefunctions for the both cases show that the particles are localized at the boundary more strongly and when $\gamma$ is bigger.
IV.3 2. Boundary Mode Band
In the case where the boundary $S-$matrix $K$ is singular, the bound mode exists and the singular term can be written as
$2\cos k-2i\gamma-J^{2}e^{ik}=0,$ (S.50)
@@ IV.2 Roots with Finite Imaginary Part
$e^{ik}=\frac{\pm\sqrt{J^{2}-1-\gamma^{2}}-i\gamma}{J^{2}-1}:=e^{-\phi},$ (S.52)
under the assumption $|e^{ik}|>1$. Plugging in $k$ into the other equation gives the solution of $q$ as
$-\frac{2\cos q-2i\gamma-J^{2}e^{iq}}{2\cos q-2i\gamma-J^{2}e^{-iq}}\left(\frac{\sin k-\sin q-2\gamma}{\sin k-\sin q+2\gamma}\right)\left(\frac{\sin k+\sin q-2\gamma}{\sin k+\sin q+2\gamma}\right)=e^{i2(L+1)q}.$ (S.53)
@@
IV.4 3. Diffusive Band
In the $L\to\infty$ limit, without lack of generality, we choose $\Im k_{1}<0$. Then the bulk scattering matrix being singular indicates the solutions satisfy the string hypothesis in the way $k_{2}=k_{1}^{*}+\pi$. These solutions form a band of solutions whose corresponding eigenvalues have a vanishing real part. Now we define $k=\Re k+i\Im k$, then the S-matrix being singular also gives the relationship between $\Re k$ and $\Im k$ as
$\Re k=\pi-\arcsin\frac{\gamma}{\cosh\Im k}.$ (S.54)
@@
$\left(\frac{2\cos\Re k\cosh\Im k-i2\sin\Re k\sinh\Im k-2i\gamma-J^{2}e^{i(\Re k-\Im k)}}{2\cos\Re k\cosh\Im k-i2\sin\Re k\sinh\Im k-2i\gamma-J^{2}e^{-i(\Re k-\Im k)}}\right)^{2}\left(\frac{2i\sinh\Im k\cos\Re k+2\gamma}{2i\sinh\Im k\cos\Re k-2\gamma}\right)^{2}=e^{-i4(L+1)\Re k}.$ (S.55)
@@
$2\arcsin(z_{n})+\mathcal{O}(L^{-1})=-\pi\frac{n}{L}$ (S.56)
with $n=0,1,2,...,L-2$. This equation which look extremely simple gives the spectrum $E_{n}=4i\sqrt{\gamma^{2}-\sin\left(\frac{n\pi}{2L}\right)^{2}}-4i\gamma+\mathcal{O}(L^{3})$, which match the numerical solution extremely well as shown in Fig. S.3. This figure shows that the analytical calculation fits the exact diagonalization reasonably well when $\gamma>1$ which the system is experiencing a true diffusive bulk. However, when $\gamma<1$, the diffusive band mixes with the free band and results in a worse but reasonable prediction because we are not taking care about the contribution from the impurity.
Noticing $\Im k<0$ gives the constraint $\Im E>-4\gamma$, and all solutions beyond this limit are not physical and should be dropped. Since $\cosh\Im k>1$, the largest imaginary energy is when $\Im k=0$, which gives $\Re k=\pi-\arcsin\gamma$ and thus resulting the energy $E_{max}=4i\sqrt{\gamma^{2}-1}-4i\gamma$. This is the lower bound of the dissipation band, and when $|E_{max}|>4\gamma$ which can only happen when $\gamma<1$ the dissipation band mixes with the free band.
IV.5 4. Double-Boundary Modes
As discussed in the main text, there is another solution when both the boundary S-matrix and the bulk S-matrix are singular. Such a mode physically describes two particles (in the imaginary Hubbard model) that are both exponentially localized at the of the chain and they experience a strong imaginary Hubbard interaction. The existence of these two states is discussed in the main text with a phase diagram in Fig.1(d). However, we are not able to separate these two singularities, so we treat them in two limits. When $\gamma<1$, the double-bound mode is perturbatively connected to the case when both particles form a noiseless bound mode, which means that
$G_{k_{0},k_{0}}(x,y)=\mathcal{N}\frac{1}{J_{x}}e^{-ik_{0}x}\frac{1}{J_{y}}e^{-ik_{0}y}$ (S.57)
@@
$E=4\cos k_{0}-4i\gamma+4i\gamma\left[\sum_{x=1}^{L}G_{k_{0},k_{0}}(x,x)^{2}+G_{k_{0},k_{0}}(0,0)\right]+\mathcal{O}(\gamma^{2}),$ (S.58)
and when $\gamma>1$, the bound mode is more localized on the boundary (decays faster) as one can see from the single boundary bound mode. Then we can treat it by solving a two-site Hamiltonian ($L=2$) as
$h=\begin{pmatrix}0&-J&-J&0\\ -J&-2i\gamma&0&-J\\ -J&0&-2i\gamma&-J\\ 0&-J&-J&0\end{pmatrix},$ (S.59)
which has two boundary bound mode with energy $E_{\pm}=-i\gamma\pm\sqrt{4J^{2}-\gamma^{2}}$. It is a double-bound mode only when $\Re E\neq 0$ which gives the boundary between the Kondo phase and the bound mode phase at the large limit $\gamma$ as $J=\gamma/2$.
V N-particle Bethe Ansatz

### v2

@@ Appendix C Noiseless XX-impurity Quench
$\frac{(J^{2}-1)e^{ik}-e^{-ik}}{(J^{2}-1)e^{-ik}-e^{ik}}=e^{i2(L+1)k}.$ (42)
There exists a bound mode with $\Im k>0$, described by $B=0$ such that the wave function is normalizable in the thermodynamic limit, i.e. $(J^{2}-1)e^{ik}-e^{-ik}=0$ which gives $k_{BM}=\frac{i}{2}\log(J^{2}-1)$ with bound mode energy $E=2\cos k=\frac{J^{2}}{\sqrt{J^{2}-1}}$ and $\Im k>0$ gives $J>\sqrt{2}$. Likewise, $A=0$ gives $k_{BM}=-\frac{i}{2}\log(J^{2}-1)$. After studying the spectrum and equilibrium physics, we can now study the quench dynamics of this problem. Since the model is non-interacting, we need to solve only the single-mode evolution and take the algebraic sum of each mode at the end of the calculation. We prepare the initial state as $F(j,t=0)=F_{0}(j)=\delta_{j,j_{0}}$, and we solve for $F(j,t)$. The dynamics of a wave function is given by the following equation
$\begin{split}F(x,t)&=\sum_{\ell,k}F^{*}_{k}(\ell)F_{0}(\ell)F_{k}(j)e^{i\epsilon_{k}t}\\ &=\sum_{k}F^{*}_{k}(j_{0})F_{k}(j)e^{i\epsilon_{k}t}.\end{split}$ (43)
Let us look at the simplest case where we prepare the fermion (spin flip) at the impurity site initial $j_{0}=0$ and ask how the fermion number (magnetization) changes as time evolves; in the Kondo phase $J<\sqrt{2}$
$n_{0}(t)=F(0,t)^{2}=\left(\int\frac{dk}{\pi}\frac{J^{2}\sin^{2}(k)}{J^{4}-4(J^{2}-1)\cos^{2}(k)}e^{-i2t\cos k}\right)^{2},$ (44)
@@
$F_{Kondo}(0,t)=\frac{1}{2J^{2}\pi}\int_{-2}^{2}d\epsilon\frac{\sqrt{4-\epsilon^{2}}}{1-\mathrm{Sign}(J^{2}-1)t_{K}^{2}\epsilon^{2}}e^{-i\epsilon t},$ (45)
where the Kondo time $t_{K}=\frac{\sqrt{|J^{2}-1|}}{J^{2}}$ is the only scale in this integral. We verify this result numerically as shown in Fig.5. In the Bound Mode Phase, the same quantity can be expressed as the following
$F_{BM}(0,t)=\frac{1}{2J^{2}\pi}\int_{-2}^{2}d\epsilon\frac{\sqrt{4-\epsilon^{2}}}{1-\mathrm{Sign}(J^{2}-1)t_{K}^{2}\epsilon^{2}}e^{-i\epsilon t}+\frac{J^{2}-2}{J^{2}-1}\cos(E_{b}t).$ (46)
Figure 5: Dynamics of the spin at the impurity site with initial Néel state. The system is in bound mode phase $J=3$ (left panel) where the analytic curve of the time-dependent spin profile is given by Eq. (49) and numerical cure is calculated for $L=40$ with exact diagonalization. The system is in the Kondo phase with $J=0.5$ (right panel) where the analytic curve of the time-dependent spin profile is given by Eq. (49) and numerical cure is calculated for $L=40$ with exact diagonalization.
$F_{K}(j,\ell,t)=\int_{0}^{\pi}\frac{dk}{2\pi}\frac{1}{J_{\ell}J_{j}}(e^{-ikj}+S_{b}(k)e^{ikj})(e^{ik\ell}+S_{b}(k)^{*}e^{-ik\ell})e^{-i2t\cos k}.$ (47)
In the bound mode phase,
$F_{B}(j,\ell,t)=F_{K}(j,\ell,t)+\frac{J^{2}}{J_{j}J_{\ell}}\frac{J^{2}-2}{J^{2}-1}\left[\frac{e^{-iE_{B}t}}{(\sqrt{J^{2}-1})^{j+\ell}}+\frac{e^{iE_{B}t}}{(-\sqrt{J^{2}-1})^{j+\ell}}\right].$ (48)
$S^{z}(\ell,t)=2\sum_{j\in\{j_{x}\}}|F(j,\ell,t)|^{2}-1.$ (49)
with $F=F_{K}$ or $F=F_{B}$ depending on which phase the system is in.
Appendix D Liouvillian Spectrum
@@
The time-independent Schrödinger equation read from Eq.(38) is
$EG(x,y)=\left[J_{x}\Delta_{x}+J_{y}\Delta_{y}+4i\gamma\delta_{x,y}+2i\gamma(\delta_{x}+\delta_{y}-2)\right]G(x,y),$ (50)
@@
$G_{k,q}(x,y)=\frac{1}{\mathcal{N}}\mathcal{S}\sum_{\alpha_{1},\alpha_{2}=\pm 1}e^{i\alpha_{1}kx+i\alpha_{2}qy}\left[A(\alpha_{1}k,\alpha_{2}q)\theta(x-y)+B(\alpha_{2}q,\alpha_{1}k)\theta(y-x)\right],$ (51)
where $\mathcal{N}$ is the normalization factor such that $1=\sum_{x,y}|G_{k,q}(x,y)|^{2}$ and $\mathcal{S}$ is the symmetrizer. Note if the transformation in Eq.(34) is not performed, then there is a factor $(-1)^{y}$ multiplying the wavefunction shown in the main text. The amplitudes are related as follows:
$\begin{split}A(-k,q)&=K(k)A(k,q),\\ A(k,-q)&=K(q)A(k,q),\\ A(k,q)&=S(k,q)B(k,q),\end{split}$ (52)
@@
$K(k)=-\frac{e^{ik}+e^{-ik}-2i\gamma-J^{2}e^{ik}}{e^{ik}+e^{-ik}-2i\gamma-J^{2}e^{-ik}},$ (53)
@@
$S(k,q)=\frac{\sin k-\sin q+2\gamma}{\sin k-\sin q-2\gamma}.$ (54)
@@
$e^{2ik(L+1)}$ $=K(k)S(k,q)S(-k,q)$ (55)
$e^{2iq(L+1)}$ $=K(q)S(q,k)S(-q,k),$ (56)
which, when written explicitly, takes the form
$\begin{split}-\frac{2\cos k-2i\gamma-J^{2}e^{ik}}{2\cos k-2i\gamma-J^{2}e^{-ik}}\left(\frac{\sin k-\sin q+2\gamma}{\sin k-\sin q-2\gamma}\right)\left(\frac{\sin k+\sin q-2\gamma}{\sin k+\sin q+2\gamma}\right)&=e^{i2(L+1)k}\\ -\frac{2\cos q-2i\gamma-J^{2}e^{iq}}{2\cos q-2i\gamma-J^{2}e^{-iq}}\left(\frac{\sin k-\sin q-2\gamma}{\sin k-\sin q+2\gamma}\right)\left(\frac{\sin k+\sin q-2\gamma}{\sin k+\sin q+2\gamma}\right)&=e^{i2(L+1)q}.\end{split}$ (57)
@@
D.1 1. The free band (Roots with vanishing Imaginary part)
By taking the norm of Eq. (57), it is easy to see that
$k_{I}=\frac{1}{L+1}\ln|K(k)S(k,q)S(-k,q)|$ (58)
@@
D.2 Roots with Finite Imaginary Part
As mentioned earlier, when the scattering matrix is singular, the Bethe roots have a non-vanishing imaginary part in the thermodynamic limit. Now, let us consider different situations with different parts of the S-matrix being singular.
Figure 6: The upper two panels show the Liouvillian spectrum (top-left) and amplitude for the wavefunction in the double-bound mode (top-right) for $J=3$, $\gamma=0.5$ and $L=30$. The perturbative calculation matches the exact diagonalization result better since $\gamma$ is small. The lower two panels show the Liouvillian spectrum (bottom-left) and amplitude for the wavefunction in the double-bound mode (bottom-right) for $J=3$, $\gamma=2$ and $L=30$. The two-site Hamiltonian calculation matches the exact diagonalization result better since $\gamma$ is big. Also, wavefunctions for both cases show that the particles are localized at the boundary more strongly and when $\gamma$ is bigger.
D.3 2. Boundary Mode Band
In the case where the boundary $S$-matrix $K$ is singular, the bound mode exists, and the singular term can be written as
$2\cos k-2i\gamma-J^{2}e^{ik}=0,$ (59)
@@ D.2 Roots with Finite Imaginary Part
$e^{ik}=\frac{\pm\sqrt{J^{2}-1-\gamma^{2}}-i\gamma}{J^{2}-1}:=e^{-\phi},$ (61)
under the assumption $|e^{ik}|>1$. Plugging $k$ into the other equation gives the solution of $q$ as
$\frac{2\cos q-2i\gamma-J^{2}e^{iq}}{2\cos q-2i\gamma-J^{2}e^{-iq}}\left(\frac{\sin k-\sin q-2\gamma}{\sin k-\sin q+2\gamma}\right)\left(\frac{\sin k+\sin q-2\gamma}{\sin k+\sin q+2\gamma}\right)=e^{i2(L+1)q}.$ (62)
@@
D.4 3. Diffusive Band
In the $L\to\infty$ limit, without lack of generality, we choose $\Im k_{1}<0$. Then, the bulk scattering matrix being singular indicates the solutions satisfy the string hypothesis in the way $k_{2}=k_{1}^{*}+\pi$. These solutions form a band of solutions whose corresponding eigenvalues have a vanishing real part. Now we define $k=\Re k+i\Im k$, then the S-matrix being singular also gives the relationship between $\Re k$ and $\Im k$ as
$\Re k=\pi-\arcsin\frac{\gamma}{\cosh\Im k}.$ (63)
@@
$\left(\frac{2\cos\Re k\cosh\Im k-i2\sin\Re k\sinh\Im k-2i\gamma-J^{2}e^{i(\Re k-\Im k)}}{2\cos\Re k\cosh\Im k-i2\sin\Re k\sinh\Im k-2i\gamma-J^{2}e^{-i(\Re k-\Im k)}}\right)^{2}\left(\frac{2i\sinh\Im k\cos\Re k+2\gamma}{2i\sinh\Im k\cos\Re k-2\gamma}\right)^{2}=e^{-i4(L+1)\Re k}.$ (64)
@@
$2\arcsin(z_{n})+\mathcal{O}(L^{-1})=-\pi\frac{n}{L}$ (65)
with n=0,1,2,,L2n=0,1,2,...,L-2. This equation which look extremely simple gives the spectrum $E_{n}=4i\sqrt{\gamma^{2}-\sin\left(\frac{n\pi}{2L}\right)^{2}}-4i\gamma+\mathcal{O}(L^{3})$, which match the numerical solution extremely well as shown in Fig. 6. This figure shows that the analytical calculation fits the exact diagonalization reasonably well when $\gamma>1$, which the system is experiencing a true diffusive bulk. However, when $\gamma<1$, the diffusive band mixes with the free band and results in a worse but reasonable prediction because we are not taking care of the contribution from the impurity.
Noticing $\Im k<0$ gives the constraint $\Im E>-4\gamma$, and all solutions beyond this limit are not physical and should be dropped. Since $\cosh\Im k>1$, the largest imaginary energy is when $\Im k=0$, which gives $\Re k=\pi-\arcsin\gamma$ and thus resulting the energy $E_{max}=4i\sqrt{\gamma^{2}-1}-4i\gamma$. This is the lower bound of the dissipation band, and when $|E_{max}|>4\gamma$, which can only happen when $\gamma<1$, the dissipation band mixes with the free band.
D.5 4. Double-Boundary Modes
As discussed in the main text, there is another solution when both the boundary S-matrix and the bulk S-matrix are singular. Such a mode physically describes two particles (in the imaginary Hubbard model) that are both exponentially localized at the end of the chain, and they experience a strong imaginary Hubbard interaction. The existence of these two states is discussed in the main text with a phase diagram in Fig.2(d). However, we are not able to separate these two singularities, so we treat them in two limits. When $\gamma<1$, the double-bound mode is perturbatively connected to the case when both particles form a noiseless bound mode, which means that
$G_{k_{0},k_{0}}(x,y)=\mathcal{N}\frac{1}{J_{x}}e^{-ik_{0}x}\frac{1}{J_{y}}e^{-ik_{0}y}$ (66)
@@
$E=4\cos k_{0}-4i\gamma+4i\gamma\left[\sum_{x=1}^{L}G_{k_{0},k_{0}}(x,x)^{2}+G_{k_{0},k_{0}}(0,0)\right]+\mathcal{O}(\gamma^{2}),$ (67)
and when $\gamma>1$, the bound mode is more localized on the boundary (decays faster), as one can see from the single boundary bound mode. Then, we can treat it by solving a two-site Hamiltonian ($L=2$) as
$h=\begin{pmatrix}0&-J&-J&0\\ -J&-2i\gamma&0&-J\\ -J&0&-2i\gamma&-J\\ 0&-J&-J&0\end{pmatrix}$, (68)
which has two boundary bound mode with energy $E_{\pm}=-i\gamma\pm\sqrt{4J^{2}-\gamma^{2}}$. It is a double-bound mode only when $\Re E\neq 0$, which gives the boundary between the Kondo phase and the bound mode phase at the large limit $\gamma$ as $J=\gamma/2$.
Appendix E N-particle Bethe Ansatz